OpenStudy (anonymous):

ruth spends 10.00 of her allowance each week on school lunches. this represents 40% of her weekly allowence. how much allowence does ruth reciecve

OpenStudy (tkhunny):

10.00 is 40% of what? 10.00 = (40%)* What? What = 10.00/(40%) = 10.00 / 0.40